OpenStudy (anonymous):

Help me solve this 2(x + 2) − (x + 1) = 8

OpenStudy (anonymous):

i am confused a bit with this equation and i would like someone to explain to me how to solve this

OpenStudy (anonymous):

first you need to distribute the 2

OpenStudy (anonymous):

then distribute a -1 for the - sign in front of the other. then, combine like terms, isolate your variable and solve.

OpenStudy (anonymous):

so 2(x+2) = 2(x) + 2(2) for the first part

OpenStudy (anonymous):

the second you are going to do -1(x) and -1(1)

OpenStudy (anonymous):

so for the first part you should have 2x+4 right?

OpenStudy (anonymous):

so for the 2nd part ou should have -x - 1. put that all together and you have 2x+2-x-1=8

OpenStudy (anonymous):

now simplify by combining like terms .. thus x+1 = 8. then you need to move that 1 over to the other side by subtracting it from each side. so x+1-1=8-1 .. now you have x = 8-1

OpenStudy (anonymous):

but wjere did u get the -1 from

OpenStudy (anonymous):

im sorry.. it should be + 1

OpenStudy (anonymous):

o nevermind i got it because of the - sign infront before the x + 1 rifght?

OpenStudy (anonymous):

o wat?

OpenStudy (anonymous):

yes it should be x = 8+1 since you have to add 1 to both sides.. my mistake

OpenStudy (anonymous):

so here you are looking for x!

OpenStudy (anonymous):

first multiply out the 2 with the stuff thats inside the brackets.

OpenStudy (anonymous):

and yes.. because of the minus sign. u distribute that as well

OpenStudy (anonymous):

2x+4-x-1=8

OpenStudy (anonymous):

good luck

OpenStudy (anonymous):

now you subtract! 2x-x+4-1=8

OpenStudy (anonymous):

here its x+3=8 not subtracy the 3 on both sides. you will get x=8-3 so the x=5

OpenStudy (anonymous):

thank you both!!!!

OpenStudy (anonymous):

u r weclome :)